# Evidence-based multimodal fusion on structured EHRs and free-text notes for ICU outcome prediction

### v1

@@ ABSTRACT
Objective: Accurate Intensive Care Unit (ICU) outcome prediction is critical for improving patient treatment quality and ICU resource allocation. Existing research mainly focuses on structured data and lacks effective frameworks to integrate clinical notes from heterogeneous electronic health records (EHRs). This study aims to explore a multimodal framework based on evidence theory that can effectively combine heterogeneous structured EHRs and free-text notes for accurate and reliable ICU outcome prediction.
Materials and Methods: We proposed an evidence-based multimodal fusion framework to predict ICU outcomes, including mortality and prolonged length of stay (PLOS), by utilizing both structured EHR data and free-text notes from the MIMIC-III database. We compare the performance against baseline models that use only structured EHRs, free-text notes, or existing multimodal approaches.
Results: The results demonstrate that the evidence-based multimodal fusion model achieved both accurate and reliable prediction. Specifically, it outperformed the best baseline by 1.05%/1.02% in BACC, 9.74%/6.04% in F1 score, 1.28%/0.9% in AUROC, and 6.21%/2.68% in AUPRC for predicting mortality and PLOS, respectively. Additionally, it improved the reliability of the predictions with a 26.8%/15.1% reduction in the Brier score and a 25.0%/13.3% reduction in negative log-likelihood.
Conclusion: This study demonstrates that the evidence-based multimodal fusion framework can serve as a strong baseline for predictions using structured EHRs and free-text notes. It effectively reduces false positives, which can help improve the allocation of medical resources in the ICU. This framework can be further applied to analyze multimodal EHRs for other clinical tasks.
@@
Keywords: Multimodal learning, Evidence fusion, Dempster-Shafer theory, ICU outcome prediction, Electronic health records
INTRODUCTION
@@
The Intensive Care Unit (ICU) is a specialized hospital ward that offers comprehensive and continuous care to critically ill patients. As the population of critically ill patients grows, the demand for ICUs has risen significantly, placing strain on already limited and costly intensive care resources 1, especially during public health crises like the COVID-19 pandemic, when hospitals face an overwhelming surge of patients.
@@
Due to the limited availability of intensive care resources, researchers have emphasized the necessity of predicting ICU outcomes such as mortality rates and prolonged lengths of stay (PLOS). Accurate predictions can help in the efficient allocation of medical resources for patients in need and reduce unnecessary expenses without compromising patient care. Furthermore, they are crucial for healthcare providers in making informed decisions about patient care strategies and providing early interventions to patients at high risk of adverse outcomes 2, 3.
@@
Over the past two decades, the adoption of electronic ICU technology has enabled the collection of extensive data on ICU patients, creating new opportunities for developing advanced methods to predict ICU outcomes. Most previous research has concentrated on modeling ICU outcome predictions using structured EHR data 4, 5, 6, which often captures only a portion of clinical information. It may miss out on the rich contextual information that unstructured EHR data (such as nursing notes, patient narratives, and imaging reports) can provide. Natural language processing (NLP) techniques have been well explored to extract valuable insights from unstructured free-text EHR data 7, 8, 9. Therefore, effective multimodal learning algorithms are essential to integrate heterogeneous EHRs for better ICU outcome prediction. Recently, deep learning-based multimodal models have been proven to combine structured EHR data and free-text data at the deep feature level for clinical outcome predictions 10, 11, 12. While those methods have increased the prediction accuracy, the question regarding the reliability of those models in noisy and unstable clinical situations still exists.
@@
In this study, we hypothesize that an effective multimodal deep neural network that incorporates both structured EHR data and free-text notes can improve the prediction of ICU outcomes. Our objective is to propose and evaluate a novel evidence-based multimodal framework for ICU outcome prediction using multimodal EHR data. Unlike traditional methods that rely on data or feature embeddings, our framework integrates information based on evidence derived from different modalities. Specifically, we use state-of-the-art deep neural networks for single-modality feature extraction: ResNet/Transformer-based models for structured EHR data and pre-trained language models for free-text EHR data. The extracted features are independently mapped into evidence and then combined in the evidence space. Experimental results on the MIMIC-III database for mortality and prolonged length of stay (PLOS) predictions demonstrate the effectiveness of the proposed model. The code implementation is available on GitHub repository (https://github.com/yuchengruan/evid_multimodal_ehr)
MATERIALS AND METHODS
Overview
In this section, we demonstrate the cohort, patient features, prediction outcomes, and proposed evidence-based multimodal fusion framework in combining structured EHRs and free-text notes for ICU outcome prediction.
Cohort
@@
MIMIC-III (Medical Information Mart for Intensive Care III) is a large, publicly available database containing de-identified health records from patients in critical care units at Beth Israel Deaconess Medical Center (from the US) between 2001 and 2012 13. We collected structured EHR data and free-text clinical notes from the database. Patients were excluded if they (1) were under 18 years of age at admission and (2) had incomplete length of stay or mortality data. For patients with multiple ICU stays, we considered only the first.
Patient features
@@
Patient features include both structured EHR data and unstructured free-text notes. In this section, we demonstrate the patient features in our study and provide details about the data preprocessing steps.
@@
Free-text notes contain a rich repository of clinical information about observations, assessments, and the overall clinical picture, which structured data often fails to capture. Furthermore, they provide an important context for interpreting structured data. For instance, while lab results may indicate abnormal values, free-text notes can clarify the relevance of these results by considering the patient’s history, comorbidities, or specific circumstances at the time of testing. As a result, NLP techniques, especially pre-trained LLM, can be applied to these notes to gain deeper insights for data-driven predictions. In this study, we focus on Nursing, Nursing/Other, Physician, and Radiology notes, as these comprise the majority of clinical documentation and are frequently recorded in the MIMIC-III database 10. We extracted only the first 24 hours of notes for each admission to facilitate early outcome prediction.
@@
Prediction Outcomes
In this study, we focus on two ICU prediction outcomes: mortality and prolonged length of stay.
@@
Length of stay refers to the duration between a patient’s admission to and discharge from the ICU. In this study, we aim to predict prolonged length of stay (PLOS), defined as a stay exceeding 7 days 14, 15, 10. Prolonged ICU stays are often linked to severe illnesses, complications, and increased mortality. Moreover, they place considerable pressure on hospital resources by reducing the availability of ICU beds and specialized personnel. Efficient management of ICU LOS not only improves patient outcomes but also enhances the overall effectiveness of healthcare systems. This problem is framed as a binary classification task.
Model architecture
The overview of the proposed framework is illustrated in Figure 1. In general, we propose to generate the modality-level evidence for both structured data and free-text notes using the evidence mapping module and fuse the modality-level evidence for final prediction.
Evidence mapping (EM)
Denœux 16 proposed an evidential neural network (ENN) that maps imperfect (uncertain, imprecise, or noise) input features into degrees of belief and ignorance (uncertainty) under the framework of Dempster-Shafer (DS) theory 17, 18. The essential concept of ENN is to consider each prototype as a piece of evidence, which is discounted based on its distance from the input vector. The evidence from different prototypes is then aggregated using Dempster’s combination rule. Inspired by ENN and its promising adoptions in medical data analysis 19, 20, 21, 22, we propose incorporating ENN as an evidence mapping module with the state-of-the-art encoders to generate evidence for structured EHRs and free-text notes. Further details about DS theory and evidence mapping can be found in Appendix A and B. Given modality level input, the evidence mapping module can output the evidence for each class as well as the uncertainty regarding this prediction.
Structured data evidence mapping
To produce modality evidence for structured data, we initially used a structured data encoder to extract deep features (the output dimension is set to 32). We considered three popular encoders to extract the embeddings: MLP, ResNet, FT-Transformer 23 (see Baselines section for more details). Subsequently, we introduced an evidence mapping module to transform the deep features into evidence embeddings for structured data.
Free-text notes evidence mapping
Similarly, we utilized pre-trained language models to extract the deep features from clinical notes, on which we developed an evidence mapping module to produce the evidence of the modality. Our primary analysis focused on pre-trained architectures similar to BERT. Accordingly, three BERT-based architectures were evaluated: BERT 24, BioBERT 24, Clinical BERT 25 (see Baselines section for more details). To minimize computational overhead while maintaining high predictive performance, we froze the pre-trained language models and fine-tuned an additional layer (128 hidden units) on top in model training.
Evidence fusion (EF)
Based on the modality-level evidence obtained from structured EHR and free-text notes, we developed the evidence fusion module based on Dempster’s combination rule to generate the final evidence for decision-making. Further details about the fusion of EHR and free-text notes evidence can be found in Appendix C.
Augmented learning algorithm
We optimize the proposed framework using an augmented learning algorithm, which includes two types of optimization objectives: (1) main objective and (2) auxiliary objective. The main objective is to optimize predictive performance based on transformed evidence as the primary loss function. Additionally, two auxiliary cross-entropy losses are incorporated to enhance the feature representation capability of the independent encoders for the two modalities, as the evidence mapping module performs more effectively with high-quality representations. More details can be found in the Appendix D.
Evaluation and experiment
Baselines
@@
To comprehensively evaluate the effectiveness of our proposed fusion framework, we employed three types of classification models as baselines: (1) models with structured data only, (2) models with free-text notes only, and (3) existing multimodal models based on both data. The configurations for the structured data models are as follows:
@@
Random Forest 26: A decision tree-based ensemble learning method for making predictions.
@@
MLP 23: A fundamental neural network with fully connected layers to encode structured data and serves as a reliable baseline. The model was configured with 3 layers, 32 hidden units, and a dropout rate of 0.1.
@@
ResNet 23: Because of the success of ResNet in computer vision 27, it has also been adapted for structured data modeling. Specifically, the main building block is simplified by providing a direct path from input to output. The configuration in our study has 3 residual blocks, 32 hidden units, and a dropout rate of 0.1.
@@
FT-Transformer 23: It converts all categorical and numerical features into embeddings, which are then processed through a couple of Transformer layers. It has demonstrated superior performance as a structured data encoder across various tasks. The model configuration has 3 Transformer layers with 192 hidden units, 8 attention heads, and a dropout rate of 0.2.
As our focus is on pre-trained language models similar to BERT, we employed three BERT-based text classification models in this study:
@@
BERT 24: a pre-trained large language model English corpus using a self-supervised approach. Through masked language modeling and next-sentence prediction objectives, the model learns internal representations of the English language, enabling it to extract features useful for downstream tasks. In this study, we utilized the google-bert/bert-base-uncased model from Huggingface 28 for feature extraction.
@@
BioBERT 29: a variant of BERT pre-trained on biomedical literature, such as PubMed abstracts, and is optimized to perform more effectively on biomedical NLP tasks. In this study, we used dmis-lab/biobert-v1.1 from Huggingface as the extraction model.
@@
Clinical BERT 25: A fine-tuned version of BERT on clinical notes from the MIMIC-III database, making it well-suited for handling medical terminology and clinical narratives to enhance performance on clinical tasks. In our study, we used emilyalsentzer/Bio_ClinicalBERT from the Huggingface transformer library for extracting embeddings from clinical notes.
@@
We also implemented a Multimodal approach for the ICU outcome prediction from 12: it concatenated the structured EHRs with the extracted embeddings from free-text notes, which were then passed into 2 fully connected layers for prediction. It supports different text encoders, and in this study, we used BERT, BioBERT, and Clinical BERT for fair comparisons
Implementation details
The dataset was randomly divided with 60% for training, 20% for validation, and 20% for testing. For model training, a mini-batch size of 32 was used, and the maximum number of epochs was set to 150, with early stopping applied. Each model was trained five times with different random seeds, and we reported the average results along with the standard error of the metrics to ensure statistical reliability. Hyperparameters were optimized for all baseline models to achieve the best results.

### v2

# Towards accurate and reliable ICU outcome prediction: a multimodal learning framework based on belief function theory using structured EHRs and free-text notes

@@ Abstract
Accurate Intensive Care Unit (ICU) outcome prediction is critical for improving patient treatment quality and ICU resource allocation. Existing research mainly focuses on structured data, e.g. demographics and vital signs, and lacks effective frameworks to integrate clinical notes from heterogeneous electronic health records (EHRs). This study aims to explore a multimodal framework based on belief function theory that can effectively fuse heterogeneous structured EHRs and free-text notes for accurate and reliable ICU outcome prediction. The fusion strategy accounts for prediction uncertainty within each modality and conflicts between multimodal data. The experiments on MIMIC-III dataset show that our framework provides more accurate and reliable predictions than existing approaches. Specifically, it outperformed the best baseline by 1.05%/1.02% in BACC, 9.74%/6.04% in F1 score, 1.28%/0.9% in AUROC, and 6.21%/2.68% in AUPRC for predicting mortality and PLOS, respectively. Additionally, it improved the reliability of the predictions with a 26.8%/15.1% reduction in the Brier score and a 25.0%/13.3% reduction in negative log-likelihood. By effectively reducing false positives, the model can aid in better allocation of medical resources in the ICU. Furthermore, the proposed method is very versatile and can be extended to analyzing multimodal EHRs for other clinical tasks. The code implementation is available on https://github.com/yuchengruan/evid_multimodal_ehr.
keywords:
@@
Multimodal Learning, Belief Function Theory, Evidence Fusion, ICU outcome prediction, Electronic Health Records
1 Introduction
@@
The Intensive Care Unit (ICU) is a specialized hospital ward that offers comprehensive and continuous care to critically ill patients. As the population of critically ill patients grows, the demand for ICUs has risen significantly, placing strain on already limited and costly intensive care resources [1], especially during public health crises like the COVID-19 pandemic, when hospitals face an overwhelming surge of patients [2].
@@
Due to the limited availability of intensive care resources, researchers have emphasized the necessity of predicting ICU outcomes such as mortality rates and prolonged lengths of stay (PLOS). Accurate predictions can help in the efficient allocation of medical resources for patients in need and reduce unnecessary expenses without compromising patient care. Furthermore, they are crucial for healthcare providers in making informed decisions about patient care strategies and providing early interventions to patients at high risk of adverse outcomes [3, 4].
@@
Over the past two decades, the adoption of electronic ICU technology has enabled the collection of extensive data on ICU patients, creating new opportunities for developing advanced methods to predict ICU outcomes. Most previous research has concentrated on modeling ICU outcome predictions using structured EHR data [5, 6, 7], which often captures only a portion of clinical information. It may miss out on the rich contextual information that unstructured EHR data (such as nursing notes, patient narratives, and imaging reports) can provide. Natural language processing (NLP) techniques have been well explored to extract valuable insights from unstructured free-text EHR data [8, 9, 10, 11]. Therefore, effective multimodal learning algorithms are essential to integrate heterogeneous EHRs for better ICU outcome prediction.
Recently, deep learning-based multimodal models have been proven to combine structured EHR data and free-text data at the deep feature level for clinical outcome predictions [12, 13, 14]. These models often simply concatenate structured data with encoded features from free texts to generate patient representations for decision-making. While those approaches have improved prediction accuracy, the clinical impacts between the two heterogeneous modalities [15] are now well explored. Another limitation of existing research is the lack of reliability evaluation for deep learning models. Unreliable predictions can lead to incorrect diagnoses or treatment plans, potentially harming patients [16, 17, 18]. Therefore, evaluating the prediction reliability is crucial beyond just predictive accuracy, especially in critical care settings. However, concerns about the reliability of existing models in noisy and unstable clinical environments still remain.
Belief function theory (BFT), also known as Dempster-Shafer theory (DST), is a powerful framework for modeling, reasoning with, and integrating imperfect (noisy, uncertain, conflicting) data [19, 20, 21]. The effectiveness of BFT in low-quality and multimodal medical image analysis has been widely reviewed in [22, 23]. However, the study of BFT in EHR data is limited. Ling et al.[24] first studied the survival prediction uncertainty using structured EHRs under the framework of BFT and possibility theory. The heterogeneity of clinical data and other medical modalities, e.g., imaging and genetic, are also studied in [25] by combining multimodal data using the generated Dempster’s combination rule [26].
@@
In this work  [^1] This work is an extended version of the short paper presented at the 8th International Conference on Belief Functions (BELIEF 2024) [27]., we further study the effectiveness of BFT in multimodal EHR analysis using structured EHRs and free-text notes with a focus on ICU outcomes prediction. We propose a multimodal learning model under the BFT framework with accurate and reliable ICU outcomes prediction using multimodal EHR data. Instead of developing more effective feature extraction or interaction strategies for multimodal data communication, our framework focuses on effective evidence fusion study and integrates information based on the evidence derived from different modalities. Specifically, we use state-of-the-art deep neural networks for single-modality feature extraction: ResNet/Transformer-based models for structured EHR data and pre-trained language models for free-text EHR data. The extracted features are independently mapped into evidence with an evidence mapping module and then combined in the evidence space in an evidence fusion module. Experimental results on the MIMIC-III database for mortality and prolonged length of stay (PLOS) predictions demonstrate the effectiveness of our proposed model in both predictive accuracy and reliability.
2 Preliminaries
2.1 Belief function theory
Belief function theory (BFT) was first introduced by Dempster and Shafer [20, 19]. The expressive capabilities of belief functions enable a more accurate representation of evidence than relying solely on probabilities. Let $\Omega=\{\omega_{1},\omega_{2},\cdots,\omega_{M}\}$ be a finite set of hypotheses about some question, called the frame of discernment. Evidence about a variable taking values in $\Omega$ can be represented by a mass function: $2^{\Omega}$ to [0,1] such that $m(\emptyset)=0$ and
$\sum_{A\subseteq\Omega}m(A)=1.$ (1)
For any hypothesis $A\subseteq\Omega$, the quantity $m(A)$ is interpreted as a share of a unit mass of belief allocated to the hypothesis that the truth is in $A$, and which cannot be allocated to any strict subset of $A$ based on the available evidence. Set $A$ is called a focal set of $m$ if $m(A)>0$. A mass function is said to be Bayesian if its focal sets are singletons, and logical if it has only one focal set. Two mass functions $m_{1}$ and $m_{2}$ representing independent items of evidence can be combined conjunctively by Dempster’s combination rule [19] $\oplus$ as
for all $A\neq\emptyset$, where $\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)$ is the degree of conflict among the two pieces of evidence, The nice information fusion attribute of BFT points out the high potential in heterogenetic medical data analysis.
After aggregating all available evidence, the final decision of BFT can be made based on the pignistic transformation proposed by Smets in the Transferable Belief Model [28] that combinese all mass functions using the following expression:
2.2 Evidential neural network
Denœux [21] proposed an evidential neural network (ENN) that maps imperfect (uncertain, imprecise, or noise) input features into degrees of belief and ignorance (uncertainty) under the framework of BFT [19, 29]. The essential concept of ENN is to consider each prototype as a piece of evidence, which is discounted based on its distance from the input vector. The evidence from different prototypes is then aggregated using Dempster’s combination rule.
As illustrated in Figure 1, the ENN consists of one input layer, one hidden layer, and one output layer. The input layer is composed of $H$ units ($H$ is the number of prototypes), whose weights vectors are prototypes $\pi_{1},\pi_{2},\cdots,\pi_{H}$ in input space. The activation of unit $h$ in the input layer is
$s_{h}=\beta_{h}\mathrm{exp}(-\gamma_{h}d_{h}^{2}),$ (4)
Figure 1: The illustration of ENN
$m_{h}(\{\omega_{c}\})$ $=u_{h}^{(c)}s_{h},\quad c=1,2,\cdots,M,$ (5a)
$m_{h}(\Omega)$ $=1-s_{h},$ (5b)
$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{H}\in\mathbb{R}^{M+1}.$ (7)
The combined mass functions (the outputs of the ENN) represent the degrees of belief about the given class with $m(\{\omega_{c}\})$, as well as its prediction uncertainty with $m(\Omega)$. In our binary classification case, the dimension of ENN outputs would be three.
3 METHODS
3.1 Model architecture
The overview of the proposed framework is illustrated in Figure 2. The general idea of this framework is to generate the modality-level evidence for both structured data and free-text notes using the evidence mapping module and fuse the modality-level evidence with Dempster’s combination rule for final prediction.
3.1.1 Evidence mapping (EM)
Inspired by ENN and its promising adoptions in medical data analysis [30, 31, 32, 33], we propose incorporating ENN as an evidence mapping module with the state-of-the-art encoders to generate evidence for structured EHRs and free-text notes. Given modality level input, the evidence mapping module can output the evidence for each class as well as the uncertainty regarding this prediction.
Structured data evidence mapping
To produce modality evidence for structured data, we initially used a structured data encoder to extract deep features (the output dimension is set to 32). We considered three popular encoders to extract the embeddings: MLP, ResNet, FT-Transformer [34] (see Baselines section for more details). Subsequently, we introduced an evidence mapping module (the number of prototypes is set to 20) to transform the deep features into evidence embeddings for structured data.
Free-text notes evidence mapping
Similarly, we utilized pre-trained language models to extract the deep features from clinical notes, on which we developed an evidence mapping module (the number of prototypes is set to 20) to produce the evidence of the modality. Our primary analysis focused on pre-trained architectures similar to BERT. Accordingly, three BERT-based architectures were evaluated: BERT [35], BioBERT [35], Clinical BERT [36] (see Baselines section for more details). To minimize computational overhead while maintaining high predictive performance, we froze the pre-trained language models and fine-tuned an additional layer (128 hidden units) on top in model training.
3.1.2 Evidence fusion (EF)
Based on the modality-level evidence obtained from structured EHR and free-text notes, we developed the evidence fusion module based on Dempster’s combination rule to generate the final evidence for decision-making.
$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{K}\in\mathbb{R}^{M+1},$ (8)
3.2 Augmented model optimization algorithm
We optimize the proposed framework using an augmented learning algorithm, which includes two types of optimization objectives: (1) main objective and (2) auxiliary objective. The main objective is to optimize predictive performance based on transformed evidence as the primary loss function. Additionally, two auxiliary cross-entropy losses are incorporated to enhance the feature representation capability of the independent encoders for the two modalities, as the evidence mapping module performs more effectively with high-quality representations.
$\mathcal{L}_{main}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}(p_{i}(\omega_{c})),$ (9)
where $\alpha,\beta$ are the hyperparameters that control the balance between the main loss and the auxiliary cross-entropy losses. In both tasks, we set $\alpha=2$ and $\beta=1$.
4 Experiments
4.1 Study Cohort
@@
This study used data from MIMIC-III (Medical Information Mart for Intensive Care III), a large, publicly available database containing de-identified health records from patients in critical care units at Beth Israel Deaconess Medical Center (from the US) between 2001 and 2012 [37]. We collected structured EHR data and free-text clinical notes from the database. Patients were excluded if they (1) were under 18 years of age at admission and (2) had incomplete length of stay or mortality data. For patients with multiple ICU stays, we considered only the first.
4.2 Input features
@@
Input features include both structured EHR data and unstructured free-text notes. In this section, we demonstrate the patient features in our study and provide details about the data preprocessing steps.
@@
Free-text notes contain a rich repository of clinical information about observations, assessments, and the overall clinical picture, which structured data often fails to capture. Furthermore, they provide an important context for interpreting structured data. For instance, while lab results may indicate abnormal values, free-text notes can clarify the relevance of these results by considering the patient’s history, comorbidities, or specific circumstances at the time of testing. As a result, NLP techniques, especially pre-trained LLM, can be applied to these notes to gain deeper insights for data-driven predictions. In this study, we focus on Nursing, Nursing/Other, Physician, and Radiology notes, as these comprise the majority of clinical documentation and are frequently recorded in the MIMIC-III database [12]. We extracted only the first 24 hours of notes for each admission to facilitate early outcome prediction.
@@
4.3 Prediction Tasks
In this study, we focus on two ICU prediction outcomes: mortality and prolonged length of stay. Since the two clinical outcomes are rare in patient popularity, we applied a simple class weighting approach during training based on relative class frequencies to mitigate biases and handle the imbalance in EHR data.
@@ 4.3 Prediction Tasks
Length of stay refers to the duration between a patient’s admission to and discharge from the ICU. In this study, we aim to predict prolonged length of stay (PLOS), defined as a stay exceeding 7 days [38, 39, 12]. Prolonged ICU stays are often linked to severe illnesses, complications, and increased mortality. Moreover, they place considerable pressure on hospital resources by reducing the availability of ICU beds and specialized personnel. Efficient management of ICU LOS not only improves patient outcomes but also enhances the overall effectiveness of healthcare systems. This problem is framed as a binary classification task.
4.4 Baselines
@@
To comprehensively evaluate the effectiveness of our proposed fusion framework, we compared it against three baseline model categories: (1) models using only structured data, (2) models using only free-text notes, and (3) existing multimodal models that integrate both data types.
Structured data baseline
The following models were used to evaluate performance with structured EHR data:
@@
  • Random Forest [40]: A decision tree-based ensemble learning method for making predictions.
@@
  • MLP [34]: A fundamental neural network with fully connected layers to encode structured data and serves as a reliable baseline. The model was configured with 3 layers, 32 hidden units, and a dropout rate of 0.1.
@@
  • ResNet [34]: Because of the success of ResNet in computer vision [41], it has also been adapted for structured data modeling. Specifically, the main building block is simplified by providing a direct path from input to output. The configuration in our study has 3 residual blocks, 32 hidden units, and a dropout rate of 0.1.
@@
  • FT-Transformer [34]: It converts all categorical and numerical features into embeddings, which are then processed through a couple of Transformer layers. It has demonstrated superior performance as a structured data encoder across various tasks. The model configuration has 3 Transformer layers with 192 hidden units, 8 attention heads, and a dropout rate of 0.2.
Free-text notes baseline
To assess text-based prediction performance, we compared our model against three BERT-based text classification approaches:
@@
  • BERT [35]: A pre-trained language model trained on a large English corpus using self-supervised learning. It learns contextual representations through masked language modeling and next-sentence prediction. In this study, we used the Google-bert/bert-base-uncased model from Huggingface [42] for feature extraction.
@@
  • BioBERT [43]: a variant of BERT pre-trained on biomedical literature, such as PubMed abstracts, and is optimized to perform more effectively on biomedical NLP tasks. In this study, we used dmis-lab/biobert-v1.1 from Huggingface as the extraction model.
@@
  • Clinical BERT [36]: A fine-tuned version of BERT on clinical notes from the MIMIC-III database, making it well-suited for handling medical terminology and clinical narratives to enhance performance on clinical tasks. In our study, we used emilyalsentzer/Bio_ClinicalBERT from the Huggingface transformer library for extracting embeddings from clinical notes.
Multimodal modal baseline
@@
To compare against multimodal models, we implemented the concatenation-based approach from [14]. This method combines structured EHR data with extracted text embeddings from free-text notes, followed by two fully connected layers for prediction. To ensure fairness, we tested this approach with BERT, BioBERT, and Clinical BERT as text encoders.
4.5 Implementation details
The dataset was randomly divided with 60% for training, 20% for validation, and 20% for testing. For model training, a mini-batch size of 32 was used, and the maximum number of epochs was set to 150, with early stopping applied. To handle data imbalance, we used a simple class weighting technique  [^2] https://scikit-learn.org/stable/modules/generated/sklearn.utils.class_weight.compute_class_weight.html based on class frequencies, as this was not the main focus of our research. The positive to negative weight ratios were set to 4.254:0.567 for the mortality prediction task and 3.660:0.579 for the PLOS prediction task.